A loan shark has offered to loan you $348,100.00 today so that you can afford the house of your dreams. The only catch is that he wants you to repay him $518,300.00 exactly 4.00 years from today. You are very tempted as you want the house really badly. What is the rate of interest on the Shark’s loan?

A couple with a new baby girl want to set aside money today that will cover her four years of college. This will be a ONE-time contribution to a mutual fund that returns 10.00% per year on average. Currently, college tuition costs $48,000.00 for four years, but it is growing at 2.00% per year. The "baby" girl will start college in 18.00 years.

How much money must they set aside today to cover the cost of college?

A couple with a new baby girl want to set aside money today that will cover her four years of college. This will be a ONE-time contribution to a mutual fund that returns 10.00% per year on average. Currently, college tuition costs $48,000.00 for four years, but it is growing at 2.00% per year. The "baby" girl will start college in 18.00 years.

What will be the cost of tuition when their "baby" girl starts college?

The average starting salary for a business school graduate is $45,850.00. With an average salary growth of 3.00% per year, how much will the graduate earn in salary 30.00 years from today?
Answer #1

Let the interest rate on loan be x

348,100 = 518,300*PVF(x%, 4 years)

PVF(x%, 4 years) = 0.6716

PVF(11%, 4 year)= 0.6587

And PVF(10%, 4 years) = 0.6830

Using interpolation, interest rate on loan = 10% + (0.6830 – 0.6716)/(0.6830 – 0.6587)

=10.47%

Fees after 18 years = 48,000*(1.02)18

=$68,555.82

Amount required today = $68,555.82*PVF(10%, 18 years)

= $68,555.82*0.180

= $12,340.05

Average Salary 30 years from now = 45,850*(1.03)30

= $111,289.98